OpenStudy (anonymous):

A 2-mi cab ride costs $4.25. A 7-mi cab ride costs $8.50. Which equation models the cost y of a cab ride that is x miles? y = 0.85x - 2.55 y = 0.85x + 2.55 y = -0.88x + 2.55 y = 0.88x - 2.55

OpenStudy (anonymous):

@skullpatrol

OpenStudy (mrnood):

This question is based on the assumption that there is a fixed charge and a variable charge for the taxi the variable charge is the 'cost per mile, and the fixed is an 'addon' regardless of distance

OpenStudy (mrnood):

First work out the difference in cost between the 2 jouneys, and aklso the diffeneence in miles between the 2 jouneys Post those numbers here and I will help you more

OpenStudy (anonymous):

I think its B but im not sure

OpenStudy (kamibug):

You can make two points from this info. (2, 4.25) and (7, 8.50) The answer choices are in slope-intercept form y = mx + b where m is the slope. So first find slope using the two points and the slope formula. \[m = \frac{ y_2 - y_1 }{ x_2 - x_1 }\] m = 8.5 - 4.25 / 7 - 2 m = 4.25 / 5 m = 0.85 Now we just need to find the y-intercept. To do that, we can use point-slope form and substitute 0 for one of the x's. point-slope form is \[y - y_1 = m(x - x_1)\] Plug in the slope and either of the points. \[y - 4.25 = 0.85(x - 2)\] Now change the other x to 0 so that we can find the y-intercept b/c the x coordinate at the y-intercept is always 0. \[y - 4.25 = 0.85(0 - 2)\] When you solve that equation for y, you should get 2.55. :) So your answer is y = 0.85 + 2.55

OpenStudy (anonymous):

thanks c:

OpenStudy (mrnood):

If you do the 2 things I wrote above, then I can help you OR you can let someone else post the complete answer and not L:EARN anything

OpenStudy (anonymous):

no I already thought it was B but she confirmed it

OpenStudy (anonymous):

yes I did to the math... but I didn't know if I did everything right..

OpenStudy (mrnood):

Well you have a good example above to follow through to check. the method :-) Good luck

OpenStudy (anonymous):

thanks I guess...